cut pint A pont R 6 is an enubi bvium pon Gwen k=44-4 cmd at vest m 0.868 by a) Determine Polen hal enengg at β eternmine ere De termine He velocity thu objest cs it passim
Answer #1

at B, Fnet = k x - m g = 0

44.4x = (0.888)(9.81)

x = 0.1962 m

(A) PE = (44.4 x 0.1962^2 / 2) - (0.88 x 9.81 x 0.1962)

PE = - 0.84 J


(B) total mechanical energy = 0


(C) Applying energy onservation,

PEi + KEi = PEf + KEf

0 = - 0.84 + 0.888 v^2 /2

v = 1.37 m/s